Stellar Company is constructing a building. Construction began on February 1 and was completed on December 31. Expenditures were $5,040,000 on March 1, $3,360,000 on June 1, and $8,400,000 on December 31.

Stellar Company borrowed $2,800,000 on March 1 on a 5-year, 12% note to help finance construction of the building. In addition, the company had outstanding all year a 8%, 5-year, $5,600,000 note payable and an 11%, 4-year, $9,800,000 note payable. Compute avoidable interest for Stellar Company. Use the weighted-average interest rate for interest capitalization purposes. (Round "Weighted-average interest rate" to 4 decimal places, e.g. 2.5125 and final answer to 0 decimal places, e.g. 5,275.)

What is Avoidable Interest?
Answer #1
Avoidable Interest $                 6,68,976
Workings:
Expenditure for the year
Mar-01 $              50,40,000 X 10 / 12 = $ 42,00,000
Jun-01 $              33,60,000 X 7 / 12 = $ 19,60,000
Dec-31 $              84,00,000 X 0 / 12 = $                -  
$           1,68,00,000 $ 61,60,000
Interest Capitalized
$                61,60,000
Less: $                28,00,000 X 12.00% = $   3,36,000
Balance $                33,60,000 X 9.91% = $   3,32,976
Interest Capitalized = $   6,68,976
Weighted Average rate of all debt:-
$              56,00,000 X 8% = $   4,48,000
$              98,00,000 X 11% = $ 10,78,000
$           1,54,00,000 $ 15,26,000
Weighted Average rate of all debt = 9.91%
($1526000 / $15400000)
Answer #2

Thank you. However, you should probably swap the WAR of all debt (WAR) and Interest Capitalized  (IC) as the WAR comes before the IC.

source: UNKWN
answered by: -A